Online Network Inc. has a net income of $510,000 in the current fiscal year. There are 100,000 shares of common stock outstanding along with convertible bonds, which have a total face value of $1.6 million. The $1.6 million is represented by 1,600 different $1,000 bonds. Each $1,000 bond pays 5 percent interest. The conversion ratio is 20. The firm is in a 30 percent tax bracket. a. Compute basic earnings per share. (Do not round intermediate calculations and round your answer to 2 decimal places.) b. Compute diluted earnings per share.
Answer #1

Basic earnings per share = Net Income/number of shares outstanding

= 510,000/100,000

= $5.1 per share

Diluted earnings per share:

Diluted earnings = 510,000+1600*1,000*5%(1-30%)

= $566,000

Dilutive shares = 100,000+1,600*20

= 132,000

Diluted earnings per share = 566,000/132,000

= $4.29 per share